2021 May

From EOVSA Wiki
Jump to navigation Jump to search

Back to Observing Log

May 16

14:14 UT The first few scans today were using the "usual" correlator equalizer coefficients (8.0). However, a test I ran earlier today indicates that my recent changes to the ADC settings has caused this equalizer setting to be quite a bit too high. After some thought, I have decided to set them to 2.0 as of this scan at 14:14 UT, which will allow most of the auto-correlations to be in a linear range. However, this will require a new expression for "desat" correction of the auto-correlation saturation. Any data taken from now may have to be reanalyzed by the pipeline, up until the date when I get the new expression defined and implemented.

13:49 UT Yesterday the new desat expressions were put into the code, so all data from now on will be properly corrected. While doing that, I discovered a bug in the SOLPNT code (introduced in late January 2021) that resulted in bad fits for the lower frequencies and hence bad calibration factors. I fixed that and checked it with the 2021 May 07 flare, and indeed the lower frequencies are far better now. Unfortunately, it means the total power calibrations should be rerun for those dates up to now (and the pipeline images and spectra that use them should also be rerun). In fact, though, that Jan 2021 change was the initial implementation of the desat correction, so really ALL of our total power calibrations should be redone. Prior to that, though, I would like to spend some time investigating the pipeline images and make sure they are the best they can be.

May 23

02:00 UT There were a lot of flares during the day on May 22. The M1 event at 17:10 UT is especially nice. We expect that the STIX instrument should have data for all of these.